« Project Status: 7 January 2008 2030UTC | Main | Project Status: 9 January 2008 0100UTC (IOP#5 Update) »

Project Status: 8 January 2008 2030UTC

IOP#5 Underway

Weather Status: The rain/ snow bands associated with the IOP#5 upper level wave arrived a bit later than expected, moving into the BLU area at 8 Jan 17UTC (11L). BLU reported heavy snow at about 19UTC (11L) and this has been continuing. Maximum precipitation rates looked to be about 0.6 inches/3 hours. Freezing levels starting at 3500-4000ft should rise a bit but BLU should remain in snow for the duration. Coarse grid models were still supporting over an inch of liquid for the event with higher resolution model still going for 1.5 to 2 inches. With multiple precip bands lined up west of the ARB it looked pretty certain that the one-inch-plus criteria would be realized. Upper trough passage is scheduled for 9 Jan 02UTC (8/18L). Precipitation showers should hang on for a few hours after trough passage. End of the event looks to be 9 Jan 06UTC (8/22L). Next wave upstream continues to favor a more northerly trajectory, basically just brushing by the ARB. Models were generally forecasting a total of 0.3 - 0.5 inches of precip, so the event continues to look weak, and not justifying an IOP. Beyond this set of disturbances the outlook is for generally dry weather with only slight shower possibilities Saturday, and nothing much beyond that. Some ensemble solutions for next week indicate a closed low off the coast that could impact the ARB if it chooses to move east.

Summary: Precipitation and IOP#5 ending between 9 Jan 03-06UTC (8/19-22L). Second wave too weak for IOP.

Observational Status: SHS and special RAOBS were initiated at 8 Jan 12UTC (04L) at the 3 sites. Radar went into operation 8 Jan 1530UTC (0730L). Crew was reporting moderate snow at BLU at telecon time. Big Bend surface observation was back on line and reported 6 inches of snow and 0.3 inches of liquid in 1.5 hours (since 8 Jan 1730UTC [0930L]). Chico profiler was back up yesterday afternoon. Cazadero was still out, as was BLU's surface site. Radar crews may have to remain on site overnight since access road has 1 foot of snow on it and at least a foot more is possible.... plowing isn't likely until early Wed. Radar crews will secure site and return to ESRL on Wed afternoon.

Decision Summary: Last sonde release 9 Jan 03UTC (19L) all sites. Radar continues until precipitation terminates sometime between 9 Jan 03-06UTC (8/19-22L)

Product Status:One HMT ensemble member had to be started manually when the system failed for that run. More products were available on the HMT web site as power is being restored at instrument sites. BLU power may be out until this weekend. ALPS workstation systems being shipped to RNO and MRY and software being transferred to SAC, so we may have ALPS for the next IOP.

McGinley,GSD

TrackBack

TrackBack URL for this entry:
http://www.etl.noaa.gov/cgi-bin/mt/mt-tb.cgi/711